### v1

# Why Chain of Thought Fails in Clinical Text Understanding

Abstract
Large language models (LLMs) are increasingly being applied to clinical care, a domain where both accuracy and transparent reasoning are critical for safe and trustworthy deployment. Chain-of-thought (CoT) prompting, which elicits step-by-step reasoning, has demonstrated improvements in performance and interpretability across a wide range of tasks. However, its effectiveness in clinical contexts remains largely unexplored, particularly in the context of electronic health records (EHRs), the primary source of clinical documentation, which are often lengthy, fragmented, and noisy. In this work, we present the first large-scale systematic study of CoT for clinical text understanding. We assess 95 advanced LLMs on 87 real-world clinical text tasks, covering 9 languages and 8 task types. Contrary to prior findings in other domains, we observe that 86.3% of models suffer consistent performance degradation in the CoT setting. More capable models remain relatively robust, while weaker ones suffer substantial declines. To better characterize these effects, we perform fine-grained analyses of reasoning length, medical concept alignment, and error profiles, leveraging both LLM-as-a-judge evaluation and clinical expert evaluation. Our results uncover systematic patterns in when and why CoT fails in clinical contexts, which highlight a critical paradox: CoT enhances interpretability but may undermine reliability in clinical text tasks. This work provides an empirical basis for clinical reasoning strategies of LLMs, highlighting the need for transparent and trustworthy approaches.
Figure 1: Performance of LLMs in clinical text understanding under Zero-shot and CoT.
1 Introduction
Large language models (LLMs) are increasingly being explored for clinical workflows, including clinical decision support, documentation drafting, information extraction, and telemedicine communication (Park et al., 2024; Vrdoljak et al., 2025). However, deploying LLMs in healthcare, an inherently safety-critical setting, requires both high accuracy and clear interpretability (Habli et al., 2020; Ennab & Mcheick, 2024). Providing explicit and transparent reasoning allows clinicians to critically evaluate LLM decision-making processes, identify gaps or misapplications of evidence, and make more informed, trustworthy decisions in patient care (Amann et al., 2020). Therefore, reliability and transparency are crucial for developing trustworthy clinical AI tools and their integration into clinical practice (Kim et al., 2025).
Chain-of-thought (CoT) prompting (Wei et al., 2022), which encourages language models to generate intermediate reasoning steps prior to producing a final answer, has become a widely adopted technique for exposing intermediate rationales. Using explicit instruction, e.g., “Let’s think step by step”, CoT prompting can unlock the latent reasoning capabilities of LLMs, substantially improving performance and perceived interpretability across various domains. This effect is especially pronounced for tasks requiring advanced reasoning and logical coherence, such as mathematical problem-solving and coding (Kojima et al., 2022).
However, the effectiveness of CoT in clinical contexts remains underexplored and poorly understood. Existing evidence largely stems from exam-style or open-domain settings (e.g., USMLE-style QA) (Wu et al., 2024a; Liévin et al., 2024) rather than real-world, multi-source clinical narratives that demand heavy factual grounding and strict differential reasoning. The clinical text in electronic health records (EHRs) is noisy, ungrammatical, and highly variable, creating additional challenges for understanding. Moreover, previous studies have revealed that LLM outputs can contain hallucinations and incorrect information, potentially introducing patient-safety risks (Asgari et al., 2025; Roustan et al., 2025). Recent analyses further caution that CoT explanations may be unfaithful, representing superficially plausible narratives that do not reflect the model’s actual decision basis, inviting over-trust if used as a reliability signal (Lyu et al., 2023). These open questions motivate a dedicated, large-scale assessment of CoT clinical text understanding.
In this study, we present the first large-scale investigation on the effectiveness of CoT prompting in LLMs in the understanding of clinical text. We evaluate 95 contemporary LLMs, including proprietary, open-source, and medical-domain variants, across 87 real-world clinical tasks spanning nine languages and eight task types, using standardized zero-shot and CoT inference protocols. Contrary to general-domain findings, we reveal that CoT consistently reduces performance despite increasing the apparent transparency of model outputs. To further investigate when and why CoT fails, we conduct four complementary analyses: (i) relating accuracy to reasoning trace length; (ii) quantifying clinical concept alignment between inputs and CoT traces; (iii) comparing lexical characteristics of CoT traces for correct vs. incorrect predictions; and (iv) auditing failure modes via an LLM-as-a-Judge with expert validation.
Overall, the main contributions of this work are:
  • A comprehensive head-to-head comparison of CoT vs. zero-shot across 95 LLMs and 87 multilingual clinical tasks, covering various clinical NLP applications under a unified evaluation protocol.
  • Empirical evidence demonstrating that CoT systematically undermines accuracy in clinical text understanding, with the larger drops for models of lower capability.
  • Mechanistic analyses linking CoT failures to longer reasoning chains and weaker clinical-concept grounding, along with characterization of lexical signatures associated with incorrect answers (e.g., clinical abbreviations and numeric mentions).
  • Development of an error taxonomy (hallucination, omission, incompleteness) evaluated via an LLM-as-a-Judge framework with clinician review, offering actionable guidance for safer use of CoT in the clinical context.
2 Related Work
2.1 Large Language Models and Chain-of-Thought
The large-scale pre-training and extensive fine-tuning equip LLM with rich world knowledge and logical reasoning capabilities, enabling them to generalize to broad tasks (Raiaan et al., 2024). Wei et al. (2022) demonstrated that CoT prompting can guide models to generate intermediate reasoning steps and significantly improve performance on tasks spanning arithmetic, commonsense, and symbolic reasoning. Then, a simple “Let’s think step by step” instruction can elicit zero-shot reasoning Kojima et al. (2022), and self-consistency Wang et al. (2023) further boosts performance by aggregating diverse reasoning paths. These techniques have become a common practice for providing detailed explanations and improving performance. Subsequently, more advanced frameworks are built on CoT: ReAct interleaves reasoning with tool use (Yao et al., 2023b); Tree-of-Thoughts (Yao et al., 2023a) and Graph-of-Thoughts Besta et al. (2024) integrate structured planning and search with CoT; Least-to-Most (Zhou et al., 2023) decomposes complex problems into solvable subproblems. Furthermore, reinforcement-learning-based reasoning models (e.g., OpenAI o1 (Jaech et al., 2024) and DeepSeek-R1 (Guo et al., )) report additional gains by training models to “think before answering,” though these advances are demonstrated mainly on math, coding, or synthetic reasoning benchmarks rather than safety-critical clinical text.
2.2 Medical Reasoning
The clinical applications of LLMs have rapidly emerged. Notably, Med-PaLM-2 (Singhal et al., 2025), integrating base-model improvements with prompting/ensemble strategies inspired by CoT, reports expert-level performance across several medical examinations. Meanwhile, In-Context Padding (Wu et al., 2024b) infers important medical entities from the input and uses them to steer the model’s reasoning trajectory, improving alignment with clinical decision pathways. CLINICR (Nachane et al., 2024) mirrors the prospective process of incremental reasoning to support differential diagnosis. These methods aim to make intermediate steps clinically grounded rather than purely free-form rationales. However, Wu et al. (2024a) and Liévin et al. (2024) also revealed that CoT does not consistently improve the performance of general LLMs on medical questions. But there is still not a systematic, large-scale evaluation of CoT on real-world clinical text tasks beyond exam-style QA. Furthermore, safety-oriented assessments in clinical documentation and summarization report hallucinations (Shah, 2024), such as fabricated facts, incorrect dosages, misattributed conditions, and instruction-following failures with potential patient-safety implications, underscoring the need for domain-specific evaluation protocols that go beyond exam-style QA and measure factual grounding in real-world clinical text (Asgari et al., 2025).
3 Methodology
3.1 Task Formulation
Let $\mathcal{T}$ denote a collection of clinical text tasks. For each task $t\in\mathcal{T}$, we consider a dataset $\mathcal{D}_{t}=\{(x_{i},y_{i})\}_{i=1}^{N_{t}}$, where $x_{i}$ is the clinical text that will serve as the model input, such as a clinical note or patient query, and $y_{i}$ is the task-specific reference output. Each task is paired with a detailed task instruction $p_{t}$ that specifies the clinical application scenario, the task requirement, and the expected output.
Given the concatenation of task instruction and input, an LLM $M$ is instructed to produce an output $\hat{y}_{i}$ according to the prompting protocol.
Specially, we investigate the model’s capabilities under two prompting strategies:
  • Zero-shot: The zero-shot prompt $p^{\text{zs}}_{t}$ is a concise, task-specific instruction that asks the model to return only the final answer in the required format, without any explanation or auxiliary text. Formally, in the zero-shot setting, the model is instructed to output only the final answer for a given input:
    $\hat{y}^{\text{zs}}_{i}\;=\;M\!\big([\,p^{\text{zs}}_{t}\,\|\,x_{i}\,]\big).$
    No intermediate reasoning is requested or scored in this mode.
  • Chain-of-thought prompting: Similar to zero-shot, the CoT prompt $p^{\text{cot}}_{t}$ includes the same task instruction and input, but adds an explicit request for step-by-step reasoning before the final answer. Formally, in the chain-of-thought (CoT) setting, the model is instructed to produce a stepwise reasoning trace before providing the final answer.
    $(r_{i},a_{i})\;=\;M\!\big([\,p^{\text{cot}}_{t}\,\|\,x_{i}\,]\big),\qquad\hat{y}^{\text{cot}}_{i}\;=\;a_{i}.$
    where $r_{i}$ is the free-form reasoning analysis generated by the model, and $a_{i}$ is the final answer; only $a_{i}$ is used for answer extraction and scoring.
Figure 2: Evaluating CoT in clinical text understanding and probing why it fails.
3.2 Clinical Text Datasets
We include $87$ clinical text datasets from the BRIDGE benchmark (Wu et al., 2025). All tasks originate from real-world clinician–patient interactions drawn from electronic health records, clinical case reports, and online consultation records, with appropriate de-identification and license compliance.
The collection spans nine languages: English, Chinese, Spanish, Japanese, German, Russian, French, Norwegian, and Portuguese. They cover 8 clinical NLP tasks to support a broad scope of clinical applications, including text classification (i.e., disease diagnosis), Named Entity Recognition (NER) and event extraction (i.e., information extraction), and Question-Answering (QA). This breadth of languages, task types, and authentic clinical settings captures the diverse reasoning demands of practice and provides a rigorous substrate for analyzing the capabilities and limitations of CoT reasoning in clinical contexts.
3.3 Models and Evaluation Protocol
We investigate $95$ widely used LLMs covering proprietary systems (e.g., Google-Gemini, Openai-GPT-4o), open-source families (e.g., Deepseek, Qwen, LLaMA, Mistral), and domain-specialized medical models (e.g., Me-LLaMA, Meditron), with parameter scales from $\sim$1B to 671B. The complete model list is provided in the Appendix Table LABEL:tab:all_model_results.
We compare performance under the two prompting strategies introduced in Sec.3.1. For each task, we leverage the task-specific scripts for label extraction and scoring provided by datasets, ensuring consistency and comparability across tasks and models.
  • For text classification, semantic similarity, natural language inference, and document-level normalization and coding, the primary metric is accuracy.
  • For NER, event extraction, and entity-level normalization and coding, we use event-level $F_{1}$ score, which requires the model to recover the target entity or event and all its associated attributes, measuring precision and recall jointly over structures.
  • For QA and summarization, we report ROUGE-average, the micro average of ROUGE-1, ROUGE-2, and ROUGE-L, to capture overlap at token and sequence levels between model response and reference.
Formally, each task $t$ specifies a deterministic scoring function $S_{t}(\hat{y},y)\in[0,1]$. Given predictions under zero-shot and CoT for the same instances, we quantify the CoT effect for a task–model pair by:
$\Delta Score\;=\;S_{t}^{\text{cot}}-S_{t}^{\text{zs}},\qquad\Delta Score(\%)\;=\;\frac{S_{t}^{\text{cot}}-S_{t}^{\text{zs}}}{S_{t}^{\text{zs}}},$
For each model $M$, the overall score is the macro-average of task-level scores across $\mathcal{T}$, yielding $\overline{S}^{\text{zs}}_{M}$ and $\overline{S}^{\text{cot}}_{M}$, and the corresponding aggregate effects $\overline{\Delta Score}_{M}$ and $\overline{\Delta Score(\%)}_{M}$.
3.4 Implementation Details
Prompt design.
To examine LLMs under broadly adopted settings, the CoT prompt follows the classical CoT instruction (Wei et al., 2022; Wang et al., 2023; Singhal et al., 2025), instantiated as brief directives like “Solve it in a step by step fashion” and then state the final answer explicitly. We do not employ any prompt optimization to avoid introducing confounding factors that could obscure the core CoT effect.
Model access and inference setting.
Details can be found in Appendix A.1.
4 Experiment and Results
To assess the practical impact of CoT in clinical contexts, we first investigate what the overall impact of CoT is on performance. Then, we analyze how CoT behaves by examining the relationship between reasoning length and performance as well as the alignment of generated content with medical concepts. Finally, we investigate why CoT fails via an error taxonomy supported by LLM-as-a-Judge assessments and clinical expert validation.
4.1 CoT Undermines LLM Performance in Clinical Text Tasks
Setup.
For each model $M$, we compute $\overline{\Delta Score}_{M}$ and $\overline{\Delta Score(\%)}_{M}$ to summarize the overall CoT effect across tasks (See 3.3). We report macro-averages at the model level and further stratify by model capability (quartiles based on $\overline{S}^{\text{zs}}_{m}$).
Figure 3: CoT prompting shows diminishing negative impact as model capability increases.
Main result and findings.
As shown in Figure 1, most LLMs ($82$ of $95$, $86.3\%$) exhibit a performance reduction under CoT prompting. Full model-wise scores are provided in Appendix Table LABEL:tab:all_model_results. Specially, even the best-performing LLMs decline under CoT: Gemini-2.5-flash ($1.6,-3.5\%$), Deepseek-R1 ($2.1,-4.9\%$), and GPT-4o ($3.5,-8.0\%$). Among the $13$ models with improvements, gains are small (0.10.12.22.2 points) and $10/13$ are below $1$ point, indicating marginal improvements.
To verify the robustness of this trend, we hand-crafted four semantically equivalent prompt pairs for each task ($p^{\text{zs}}$ and $p^{\text{cot}}$) and re-ran evaluations on three widely used open-source model families, Llama-3.3-70B-Instruct, Mistral-Large-Instruct-2411, Qwen2.5-72B-Instruct and one high-performing variant (Athene-V2-Chat). Across all variants, the direction of the effect remained consistent, and the magnitudes are were comparable (See Appendix Table S3), supporting the robustness of this global finding.
Grouping models by zero-shot capability reveals a graded pattern: higher-capability models are relatively more robust under CoT, whereas lower-capability models suffer larger drops. The linear relationship ($R^{2}=0.973$) implies a systematic interaction between model capability and the effectiveness of reasoning strategies. This stratification suggests that generating reliable reasoning imposes higher competence demands on the model.
4.2 Longer Reasoning Is Associated with Higher Performance Decline
Figure 4: Performance under CoT prompting decreases as reasoning length increases.
Setup.
We extract the reasoning trace $r_{i}$ (the text before the final answer) using the output format required by $p^{\text{cot}}_{t}$. The reasoning length $L(r_{i})$ is measured as the word count of $r_{i}$; word-level counts avoid tokenizer-specific artifacts. Within each model–task pair, instances are sorted by $L(r_{i})$ and partitioned into ten equal-sized bins (deciles). For each decile $d$, we compute the task score under CoT, $S_{t}^{\text{cot}}(d)$, and—on the same items—the zero-shot score, $S_{t}^{\text{zs}}(d)$. We then summarize the length-conditioned relative change
$\Delta Score(d)(\%)\;=\;\frac{S_{t}^{\text{cot}}(d)-S_{t}^{\text{zs}}(d)}{S_{t}^{\text{zs}}(d)}.$
Pairing by instance controls for difficulty: each decile compares CoT and zero-shot on identical inputs. We report results for four models—GPT-4o, Qwen2.5-72B-Instruct, Llama-3.3-70B-Instruct, and MeLLaMA-70B-chat, which are representative proprietary and open-source model families, and one medical variant.
Main results and findings.
Figure 4 shows a clear negative association between reasoning length and performance for all four models. Both CoT and zero-shot scores decrease as $L(r)$ increases, reflecting that longer traces often correspond to harder instances. However, the relative CoT drop grows with length: $\Delta\%(d)$ becomes increasingly negative from short to long deciles. For example, GPT-4o exhibits a widening decline of roughly 551515 percentage points across deciles, while MeLLaMA-70B-chat ranges from about $5$ to $22$ points. Zero-shot scores computed on the identical items are comparatively flat or decline much less, indicating that the amplified degradation is not merely a consequence of instance difficulty.
Figure 5: Impact of medical concept alignment on performance. Both Overlap and Coverage alignment metrics show systematic performance improvements across percentiles.
4.3 Medical Concept Alignment Mitigates CoT Degradation
Setup.
To examine how clinically grounded content relates to CoT reliability, we extract medical concepts from both the input $x_{i}$ and the CoT reasoning trace $r_{i}$ using cTAKES (Savova et al., 2010) (version 4.0.0.1), a clinical text analysis and knowledge extraction system that extracts clinical entities in free text and maps them into Concept Unique Identifiers (CUIs). For each instance, we define the set of input concepts $C_{x}$ and the set of concepts mentioned in the reasoning trace $C_{r}$ after normalization and de-duplication. We quantify alignment using two metrics to reflect the concept alignment:
$\text{Alignment-Overlap}=Jaccard(C_{r},C_{x})=\frac{|\,C_{r}\cap C_{x}\,|}{|\,C_{r}\,|\cup|C_{x}\,|},\quad\text{Alignment-Coverage}=\frac{|\,C_{r}\cap C_{x}\,|}{|\,C_{x}\,|}W$
Intuitively, Alignment-Overlap captures the extent to which the concepts introduced by the model overlap with those already present in the input, while Alignment-Coverage measures how much of the input’s clinical content is covered by the reasoning trace.
Following the above length analysis protocol, we rank instances within each model–task pair by the alignment metric and partition them into ten equal-sized bins (deciles). For each decile $d$, we compute the task score under CoT, $S_{t}^{\text{cot}}(d)$, and—on the identical instances—the zero-shot score, $S_{t}^{\text{zs}}(d)$, and report the relative change (Sec. §4.2). To capture representative model behavior while ensuring sufficient CoT data, we include models whose valid-response rate under CoT exceeds $90\%$.
Main results and findings.
Figure 5 summarizes the results. Both task performance and the CoT performance drop improve with stronger concept alignment. When Alignment-Overlap (left) increases from the lowest to the highest deciles, average scores rise steadily for both protocols, and the relative CoT deficit shrinks. A similar pattern holds for Alignment-Coverage (right): covering more of the input’s clinical concepts is associated with higher accuracy and a smaller CoT penalty. These trends indicate that CoT degradation is attenuated when the reasoning trace remains tightly grounded in the clinical content of the input, suggesting that concept-faithful rationales are a key precondition for reliable CoT.
Figure 6: Lexical contrast between CoT-correct and CoT-incorrect traces. Each point is a word; axes show normalized frequency in two groups. Color encodes density from infrequent (yellow) to frequent (blue/red). The right panels list the most characteristic words for each group.
4.4 Lexical Signatures of CoT Reasoning Trace for Correct and Incorrect Instances
Setup.
We partition the evaluation instances into two disjoint groups based on the correctness of the final answer under CoT: CoT-correct and CoT-incorrect. For each instance, we take the CoT reasoning trace $r_{i}$ and compute corpus-level word statistics after lowercasing, punctuation stripping, and lemmatization. To reduce confounding from input copying, we remove tokens that also appear in the corresponding input $x_{i}$. Let $V$ be the vocabulary after stopword removal. For a term $w\in V$, let $f_{w}^{\mathsf{cor}}$ and $f_{w}^{\mathsf{inc}}$ denote its token counts in CoT-correct and CoT-incorrect traces, with $N^{\mathsf{cor}}=\sum_{w}f_{w}^{\mathsf{cor}}$ and $N^{\mathsf{inc}}=\sum_{w}f_{w}^{\mathsf{inc}}$. We report (i) normalized frequencies $p_{w}^{\mathsf{cor}}=f_{w}^{\mathsf{cor}}/N^{\mathsf{cor}}$ and $p_{w}^{\mathsf{inc}}=f_{w}^{\mathsf{inc}}/N^{\mathsf{inc}}$ and (ii) differential usage via informative log-odds with a Dirichlet prior (Monroe et al., 2008; Kessler, 2017):
$\delta_{w}=\log\frac{f_{w}^{\mathsf{cor}}+\alpha_{w}}{N^{\mathsf{cor}}+\alpha_{0}-(f_{w}^{\mathsf{cor}}+\alpha_{w})}\;-\;\log\frac{f_{w}^{\mathsf{inc}}+\alpha_{w}}{N^{\mathsf{inc}}+\alpha_{0}-(f_{w}^{\mathsf{inc}}+\alpha_{w})},\quad z_{w}=\frac{\delta_{w}}{\sqrt{\sigma^{2}_{w}}},$
where $\alpha_{w}$ is the background frequency scaled by $\alpha_{0}$ and $\sigma^{2}_{w}$ is the approximate variance from Monroe et al. (2008). This prior-regularized contrast mitigates small-sample artifacts and yields a ranked list of words characteristic of each group.
Main results and findings.
Figure 6 contrasts term frequencies between CoT-correct and CoT-incorrect traces. Terms in the right-lower region (high in CoT-incorrect, low in CoT-correct) are disproportionately associated with errors. We observe that numbers, units, and measurement-related tokens are overrepresented among CoT-incorrect traces (e.g., pt, ct, numeric values), as are abbreviations for laboratory and physiologic assessments such as plt (platelet count) and pft (pulmonary function test). Conversely, CoT-correct traces contain a higher proportion of disease and therapy terms that are semantically aligned with the input clinical background. These lexical signatures align with our earlier analyses: tasks requiring precise handling of quantitative data and shorthand notations are more vulnerable under CoT, consistent with reports that LLMs are brittle with numeric reasoning (Akhtar et al., 2023; Khandekar et al., 2024). and domain-specific abbreviations (Rajkomar et al., 2022; Kugic et al., 2024).
(a) Distribution of error combinations
Model Hallucination (%) Omission (%) Incompleteness (%)
GPT-4o 7.7 1.7 0.2
Qwen2.5-72B 23.1 1.7 0.5
Llama-3.3-70B 25.4 3.0 0.3
Me-LLaMA-70B 27.5 87.8 96.8
(b) Prevalence of hallucination, omission, and incompleteness error types
Figure 7: Error taxonomy of CoT reasoning traces using LLM-as-a-Judge. Entries are the count/percentage of errors recognized (lower is better).
4.5 Error Analysis: Hallucination, Omission, and Incompleteness
Setup.
To characterize failure modes in the CoT reasoning trace $r$, we adopt an LLM-as-a-Judge protocol. Following prior studies about LLM-as-a-Judge on clinical task (Liu et al., 2024; Gu et al., 2024), we evaluate three primary dimensions:
  • Hallucination. Claims are unsupported by the input or medical knowledge.
  • Omission. Omission of task-relevant or clinically important facts from the input.
  • Incompleteness. Truncated, underspecified, or insufficient to justify the answer.
We adopt OpenAI-o3 as the judge model. The judge model was manually assessed by a clinical expert in a random subset ($n=200$). The full prompts can be found in Appendix A.5. To minimize bias, model identities are anonymized and responses are randomly ordered during judging. We analyze three challenging EHR-based tasks—mortality prediction on MIMIC-III-Outcome (van Aken et al., 2021), disease diagnosis on MIMIC-IV-DiReCT dataset (Wang et al., 2024) and MIMIC-IV-CDM dataset (Hager et al., 2024).
Main results and findings.
Figure 7 reports the distribution of detected error types across models. Generally, the best-performing models, GPT-4o, show the fewest errors; the following Qwen2.5-72B-Instruct and Llama-3.3-70B-Instruct are comparable. However, MeLLaMA-70B-Chat, trained primarily for medical Q&A rather than EHR cases, exhibits substantially higher omission and incompleteness rates, indicating difficulty in extracting and organizing patient-specific evidence from longitudinal notes. Overall, hallucination and omission are the predominant failure modes (Figure 7). The expert review highlights cases where historical conditions mentioned in the input are incorrectly treated as current. Compared to expert annotation, the positive predictive value (PPV) of the judge was $97.22\%$ for hallucination, $97.83\%$ for omission, and $100\%$ for incompleteness, demonstrating the reliability of the LLM-as-a-Judge.
5 Limitations
This study has several limitations. First, the scale of expert validation is limited due to time and budget constraints; although the audited sample supports the reliability of LLM-as-a-Judge, broader expert annotation can further refine the judging rubric and provide more key insights. Second, our goal is to investigate the effectiveness of CoT in general settings; we do not evaluate engineering interventions such as self-consistency, retrieval-augmented generation (RAG), or few-shot learning, which may improve model performance.
6 Conclusion
We conducted a large-scale, systematic investigation of LLM reasoning in clinical contexts, spanning 95 LLMs and real-world clinical text tasks, and show that reliable reasoning in clinical settings sets a higher demand than general domains, with current models insufficiently adapted to clinical understanding. Analyses on CoT reasoning trace implicate error accumulation in longer chains and insufficient grounding to clinical concepts as key mechanisms, exposing weaknesses in clinical abbreviations and numerical reasoning. An LLM-as-a-Judge framework with expert review also highlights the hallucination and information omission as dominant failure modes of CoT. Overall, these results establish a foundational evidence for when and why CoT underperforms in clinical settings. More efforts are needed to enhance LLMs’ understanding of medical knowledge and clinical contexts, and then develop clinically aligned reasoning strategies that reconcile transparency with trustworthy performance.
Appendix A Appendix
A.1 Implementation details
Model Access and Environment
All open-source models (except Deepseek-R1) are executed locally on NVIDIA A100 GPUs ($\times8$0 GB). Deepseek-R1 and all proprietary models are accessed through an institution-managed Azure cloud service, avoiding the potential privacy concerns and security risks.
Inference protocol and decoding controls.
All models are evaluated under a unified decoding configuration to ensure compute parity. Unless otherwise noted, we use deterministic decoding (temperature =0=0; sampling disabled). The confidence interval is estimated via bootstrap resampling with 1,000 iterations.
A.2 Model performance
We conducted a comprehensive comparison of 95 state-of-the-art large language models (LLMs) across 87 multilingual clinical tasks derived from the BRIDGE benchmark, encompassing nine languages and eight distinct task categories (classification, extraction, question answering, summarization, and others). The evaluated models comprised proprietary systems (e.g., GPT-4o and Gemini), open-source architectures (e.g., DeepSeek, Llama, and Qwen), and specialized medical domain models (e.g., Me-LLaMA and Meditron).
Our findings, presented in Table LABEL:tab:all_model_results, demonstrate that most LLMs (82 of 95, 86.3%) exhibit degraded performance under CoT inference.
Table S1: Overall Performance of LLMs Across Different Inference Strategies
Model
Model
Size
Model
Domain
Model
Percentile
Zero-shot
Score
CoT
Score
$\Delta$ Score
(CoT - ZS)
$\Delta$ (%)
(CoT - ZS)
Gemini-2.5-flash / General 80-100% 44.8 [44.1,45.6] 43.3 [42.5,44.1] 1.6-1.6 3.5-3.5
Deepseek-R1 671 General 80-100% 44.2 [43.5,45.0] 42.1 [41.3,42.9] 2.1-2.1 4.9-4.9
GPT-4o-0806 / General 80-100% 44.2 [43.4,45.0] 40.7 [39.9,41.4] 3.5-3.5 8.0-8.0
Qwen3-Next-80B-A3B-Thinking 80 General 80-100% 43.9 [43.1,44.6] 42.9 [42.1,43.7] 1.0-1.0 2.2-2.2
Gemini-1.5-pro-002 / General 80-100% 43.8 [43.1,44.6] 40.5 [39.7,41.3] 3.3-3.3 7.6-7.6
Gemini-2.0-flash-001 / General 80-100% 43.0 [42.2,43.8] 42.0 [41.2,42.8] 1.1-1.1 2.4-2.4
Mistral-Large-Instruct-2411 123 General 80-100% 42.3 [41.5,43.1] 38.9 [38.1,39.7] 3.4-3.4 8.0-8.0
Qwen3-30B-A3B-Thinking-2507 30 General 80-100% 41.8 [41.0,42.6] 41.4 [40.6,42.2] 0.4-0.4 0.9-0.9
Athene-V2-Chat 72 General 80-100% 41.7 [40.9,42.5] 39.3 [38.6,40.1] 2.4-2.4 5.6-5.6
Qwen3-235B-A22B-Thinking 235 General 80-100% 41.6 [40.8,42.5] 40.1 [39.4,40.9] 1.5-1.5 3.6-3.6
Qwen2.5-72B-Instruct 72 General 80-100% 41.6 [40.8,42.4] 38.9 [38.1,39.7] 2.8-2.8 6.6-6.6
Qwen3-32B-Thinking 32 General 80-100% 41.0 [40.3,41.8] 34.5 [33.7,35.3] 6.6-6.6 16.0-16.0
HuatuoGPT-o1-72B 72 Medical 80-100% 41.0 [40.2,41.8] 38.1 [37.4,38.9] 2.9-2.9 7.0-7.0
Qwen3-30B-A3B-Thinking 30 General 80-100% 40.9 [40.1,41.8] 39.4 [38.6,40.1] 1.6-1.6 3.9-3.9
medgemma-27b-it 27 Medical 80-100% 40.8 [40.0,41.6] 38.2 [37.5,38.9] 2.6-2.6 6.4-6.4
Qwen3-14B-Thinking 14 General 80-100% 40.2 [39.4,41.0] 37.1 [36.3,37.9] 3.1-3.1 7.7-7.7
Qwen3-4B-Thinking-2507 4 General 80-100% 40.0 [39.2,40.8] 38.8 [38.0,39.7] 1.2-1.2 3.0-3.0
Qwen3-8B-Thinking 8 General 80-100% 40.0 [39.2,40.8] 37.1 [36.2,37.9] 2.9-2.9 7.3-7.3
gemma-3-27b-it 27 General 80-100% 39.9 [39.1,40.7] 37.5 [36.8,38.3] 2.4-2.4 5.9-5.9
Qwen2.5-32B-Instruct 32 General 60-80% 39.9 [39.1,40.7] 37.7 [36.9,38.5] 2.2-2.2 5.6-5.6
Llama-3.3-70B-Instruct 70 General 60-80% 39.9 [39.1,40.7] 36.8 [36.1,37.6] 3.0-3.0 7.6-7.6
Qwen3-Next-80B-A3B-Instruct 80 General 60-80% 39.8 [39.0,40.6] 40.5 [39.7,41.3] $0.7$ $1.7$
Deepseek-R1-Distill-Llama-70B 70 General 60-80% 39.8 [39.0,40.6] 38.9 [38.2,39.7] 0.8-0.8 2.1-2.1
Deepseek-R1-Distill-Qwen-32B 32 General 60-80% 39.7 [39.0,40.5] 38.7 [37.9,39.5] 1.0-1.0 2.6-2.6
Mistral-Small-3.1-24B-Instruct-2503 24 General 60-80% 39.7 [38.9,40.6] 36.2 [35.4,37.0] 3.5-3.5 8.8-8.8
QWQ-32B 32 General 60-80% 39.4 [38.6,40.2] 37.0 [36.2,37.8] 2.3-2.3 6.0-6.0
Qwen3-32B-Non-Thinking 32 General 60-80% 39.3 [38.5,40.1] 32.8 [32.0,33.6] 6.5-6.5 16.5-16.5
QwenLong-L1-32B 32 General 60-80% 39.2 [38.4,40.1] 36.7 [35.9,37.5] 2.5-2.5 6.5-6.5
Qwen3-235B-A22B-Non-Thinking 235 General 60-80% 39.2 [38.4,40.0] 38.0 [37.2,38.8] 1.2-1.2 3.1-3.1
Llama-3.1-70B-Instruct 70 General 60-80% 39.1 [38.3,39.9] 35.1 [34.3,35.9] 4.0-4.0 10.2-10.2
Qwen3-4B-Thinking 4 General 60-80% 38.5 [37.7,39.3] 37.0 [36.2,37.8] 1.5-1.5 4.0-4.0
gemma-2-27b-it 27 General 60-80% 38.2 [37.4,39.0] 34.2 [33.4,35.0] 4.0-4.0 10.4-10.4
Baichuan-M2-32B 32 Medical 60-80% 38.0 [37.0,39.0] 38.3 [37.4,39.2] $0.3$ $0.7$
Magistral-Small-2506 24 General 60-80% 37.6 [36.8,38.4] 34.2 [33.3,35.0] 3.4-3.4 9.1-9.1
Mistral-Small-24B-Instruct-2501 24 General 60-80% 37.5 [36.7,38.4] 31.6 [30.8,32.4] 5.9-5.9 15.8-15.8
gemma-3-12b-it 12 General 60-80% 37.3 [36.6,38.1] 35.4 [34.6,36.1] 1.9-1.9 5.2-5.2
gpt-oss-120b 120 General 60-80% 37.2 [36.4,38.1] 32.1 [31.4,32.9] 5.1-5.1 13.8-13.8
Deepseek-R1-0528-Qwen3-8B 8 General 60-80% 36.9 [36.1,37.8] 36.3 [35.5,37.1] 0.7-0.7 1.8-1.8
Qwen3-30B-A3B-Instruct-2507 30 General 40-60% 36.9 [36.1,37.6] 37.5 [36.7,38.3] $0.6$ $1.7$
Qwen3-14B-Non-Thinking 14 General 40-60% 36.8 [36.0,37.6] 34.6 [33.8,35.4] 2.2-2.2 6.0-6.0
Qwen3-30B-A3B-Non-Thinking 30 General 40-60% 36.2 [35.4,36.9] 36.9 [36.1,37.7] $0.7$ $1.9$
Phi-4 14 General 40-60% 36.1 [35.4,36.9] 32.6 [31.8,33.4] 3.5-3.5 9.8-9.8
Baichuan-M1-14B-Instruct 14 Medical 40-60% 36.1 [35.3,36.9] 34.4 [33.6,35.2] 1.7-1.7 4.8-4.8
gpt-35-turbo-0125 / General 40-60% 35.3 [34.5,36.1] 31.6 [30.8,32.4] 3.7-3.7 10.4-10.4
Mistral-Small-Instruct-2409 22 General 40-60% 35.2 [34.4,36.0] 31.2 [30.4,32.0] 4.0-4.0 11.4-11.4
Llama-4-Scout-17B-16E-Instruct 109 General 40-60% 35.1 [34.4,35.9] 29.4 [28.7,30.1] 5.7-5.7 16.3-16.3
gemma-2-9b-it 9 General 40-60% 35.1 [34.3,35.9] 29.9 [29.2,30.7] 5.1-5.1 14.6-14.6
Deepseek-R1-Distill-Qwen-14B 14 General 40-60% 34.3 [33.5,35.1] 34.8 [34.0,35.6] $0.5$ $1.5$
Qwen3-4B-Instruct-2507 4 General 40-60% 34.1 [33.3,34.8] 36.2 [35.4,37.0] $2.2$ $6.3$
HuatuoGPT-o1-70B 70 Medical 40-60% 33.9 [33.2,34.6] 32.3 [31.6,33.1] 1.5-1.5 4.6-4.6
Qwen3-8B-Non-Thinking 8 General 40-60% 33.9 [33.1,34.7] 33.3 [32.5,34.1] 0.6-0.6 1.6-1.6
Llama-3-70B-UltraMedical 70 Medical 40-60% 33.4 [32.6,34.2] 29.4 [28.7,30.2] 4.0-4.0 11.9-11.9
Qwen3-4B-Non-Thinking 4 General 40-60% 33.3 [32.5,34.1] 33.2 [32.4,34.0] 0.1-0.1 0.3-0.3
Llama3-OpenBioLLM-70B 70 Medical 40-60% 33.0 [32.2,33.8] 28.8 [28.0,29.5] 4.2-4.2 12.8-12.8
Llama-3.1-Nemotron-70B-Instruct-HF 70 General 40-60% 32.7 [32.0,33.5] 24.1 [23.5,24.7] 8.7-8.7 26.4-26.4
MeLLaMA-70B-chat 70 Medical 40-60% 32.3 [31.5,33.1] 29.2 [28.4,30.1] 3.0-3.0 9.3-9.3
Yi-1.5-34B-Chat-16K 34 General 40-60% 32.1 [31.4,32.9] 29.6 [28.8,30.3] 2.5-2.5 7.9-7.9
QwQ-32B-Preview 32 General 20-40% 31.7 [31.0,32.5] 23.3 [22.5,24.1] 8.4-8.4 26.6-26.6
Qwen2.5-7B-Instruct 7 General 20-40% 31.3 [30.5,32.1] 30.2 [29.5,31.0] 1.1-1.1 3.4-3.4
Ministral-8B-Instruct-2410 8 General 20-40% 30.4 [29.5,31.2] 25.9 [25.1,26.7] 4.5-4.5 14.7-14.7
HuatuoGPT-o1-7B 7 Medical 20-40% 29.6 [28.8,30.4] 26.7 [26.0,27.4] 2.9-2.9 9.8-9.8
Phi-3.5-MoE-instruct 42 General 20-40% 29.5 [28.8,30.3] 25.3 [24.5,26.0] 4.3-4.3 14.5-14.5
medgemma-4b-it 4 Medical 20-40% 29.4 [28.3,30.5] 28.5 [27.6,29.3] 0.9-0.9 3.1-3.1
gpt-oss-20b 20 General 20-40% 29.0 [28.3,29.8] 24.9 [24.1,25.6] 4.2-4.2 14.4-14.4
Llama-3.1-8B-Instruct 8 General 20-40% 29.0 [28.2,29.8] 29.4 [28.6,30.2] $0.4$ $1.4$
Yi-1.5-9B-Chat-16K 9 General 20-40% 28.8 [28.0,29.6] 25.4 [24.6,26.2] 3.4-3.4 11.9-11.9
gemma-3-4b-it 4 General 20-40% 28.6 [27.8,29.4] 28.2 [27.4,29.0] 0.4-0.4 1.3-1.3
Deepseek-R1-Distill-Llama-8B 8 General 20-40% 28.5 [27.7,29.2] 27.3 [26.6,28.1] 1.1-1.1 4.0-4.0
Qwen2.5-3B-Instruct 3 General 20-40% 26.6 [25.8,27.3] 25.4 [24.7,26.2] 1.1-1.1 4.3-4.3
Qwen3-1.7B-Thinking 1.7 General 20-40% 26.3 [25.5,27.0] 27.7 [27.0,28.4] $1.4$ $5.4$
HuatuoGPT-o1-8B 8 Medical 20-40% 25.9 [25.1,26.6] 22.5 [21.8,23.2] 3.4-3.4 13.0-13.0
Phi-3.5-mini-instruct 4 General 20-40% 25.4 [24.7,26.1] 23.9 [23.2,24.6] 1.5-1.5 5.9-5.9
Deepseek-R1-Distill-Qwen-7B 7 General 20-40% 25.3 [24.5,26.0] 23.9 [23.1,24.7] 1.4-1.4 5.5-5.5
Llama-2-70b-chat 70 General 20-40% 25.2 [24.4,25.9] 19.0 [18.3,19.7] 6.1-6.1 24.4-24.4
Phi-4-mini-instruct 4 General 20-40% 24.5 [23.8,25.3] 22.5 [21.8,23.2] 2.0-2.0 8.3-8.3
Llama-3.2-3B-Instruct 3 General 20-40% 22.9 [22.1,23.7] 21.6 [20.9,22.3] 1.3-1.3 5.7-5.7
Qwen2.5-1.5B-Instruct 1.5 General 0-20% 22.2 [21.4,22.9] 19.5 [18.7,20.2] 2.7-2.7 12.1-12.1
Qwen3-1.7B-Non-Thinking 1.7 General 0-20% 22.0 [21.2,22.7] 22.8 [22.0,23.5] $0.8$ $3.7$
Phi-4-mini-reasoning 4 General 0-20% 21.3 [20.6,22.0] 19.8 [19.1,20.5] 1.5-1.5 7.1-7.1
Llama-2-13b-chat 13 General 0-20% 20.9 [20.2,21.7] 16.2 [15.5,16.9] 4.7-4.7 22.4-22.4
MeLLaMA-13B-chat 13 Medical 0-20% 20.8 [20.0,21.5] 20.3 [19.5,21.0] 0.5-0.5 2.4-2.4
BioMistral-7B 7 Medical 0-20% 20.4 [19.7,21.2] 10.8 [10.2,11.5] 9.6-9.6 46.9-46.9
Qwen3-0.6B-Thinking 0.6 General 0-20% 20.4 [19.6,21.1] 18.9 [18.2,19.7] 1.4-1.4 7.1-7.1
MMed-Llama-3-8B 8 Medical 0-20% 20.4 [19.7,21.1] 16.2 [15.4,16.9] 4.2-4.2 20.6-20.6
Llama-3.1-8B-UltraMedical 8 Medical 0-20% 20.2 [19.5,20.9] 18.3 [17.7,19.0] 1.8-1.8 9.0-9.0
OpenThinker3-7B 7 General 0-20% 19.8 [19.0,20.5] 18.4 [17.7,19.1] 1.4-1.4 7.2-7.2
MedReason-8B 8 Medical 0-20% 18.2 [17.4,18.9] 18.3 [17.6,19.0] $0.1$ $0.7$
Llama-2-7b-chat 7 General 0-20% 16.5 [15.8,17.2] 13.7 [13.0,14.3] 2.8-2.8 17.1-17.1
gemma-3-1b-it 1 General 0-20% 15.7 [15.0,16.4] 13.5 [12.8,14.2] 2.2-2.2 14.0-14.0
meditron-70b 70 Medical 0-20% 15.7 [15.0,16.4] 13.2 [12.5,13.9] 2.5-2.5 16.0-16.0
Qwen3-0.6B-Non-Thinking 0.6 General 0-20% 15.2 [14.5,15.9] 16.6 [15.9,17.3] $1.4$ $9.1$
Deepseek-R1-Distill-Qwen-1.5B 1.5 General 0-20% 14.3 [13.6,15.0] 13.4 [12.7,14.1] 0.8-0.8 5.9-5.9
Llama3-OpenBioLLM-8B 8 Medical 0-20% 14.2 [13.5,14.9] 13.3 [12.6,14.0] 0.9-0.9 6.4-6.4
Llama-3.2-1B-Instruct 1 General 0-20% 12.7 [12.1,13.4] 11.9 [11.2,12.5] 0.9-0.9 6.7-6.7
Phi-4-reasoning-plus 14 General 0-20% 10.6 [10.0,11.2] 10.8 [10.2,11.4] $0.3$ $2.5$
Phi-4-reasoning 14 General 0-20% 10.0 [9.4,10.6] 10.1 [9.5,10.7] $0.1$ $0.9$
meditron-7b 7 Medical 0-20% 9.5 [8.9,10.2] 9.5 [8.9,10.1] $0.0$ $0.0$
A.3 Model Performance Percentile
We defined a model’s basic capability as its macro-average zero-shot performance across all BRIDGE tasks. To examine whether model capability influences susceptibility to CoT-induced performance degradation, we stratified the 95 models into five equal percentile tiers (19 models each) to analyze this relationship systematically.
As shown in Table S2, higher-performing models are more resilient to CoT inference, with top-tier models showing smaller performance drops than lower-performing counterparts.
Table S2: Performance Drop of CoT Prompting Across LLM Strength Tiers
Model Performance Tier $\Delta$ Score (%) CoT Observation
Top 20% -5.82 Smallest gap
20%–40% -6.67 Moderate
40%–60% -7.19 Moderate
60%–80% -8.63 Large
Bottom 20% -9.86 Largest gap
A.4 Prompt design and variants
To ensure our findings were not artifacts of specific prompt formulations, we tested four additional prompt variations for both zero-shot and CoT inference using established prompting methodologies. Zero-shot variants differed in formality and constraint specification, while CoT variants used diverse reasoning framings. Complete prompt variations are shown below.
We selected 4 representative models that span different architectures and capabilities to provide a balanced assessment of prompt robustness while maintaining experimental feasibility. As shown in Table S3, the performance decrease with CoT inference was consistent across all prompt variations, confirming that our results represent a systematic limitation.
Method 1 Zero-Shot: “Return your answer in the following format. DO NOT GIVE ANY EXPLANATION:” Chain-of-Thought: “Solve it in a step-by-step fashion, return your answer in the following format, PROVIDE DETAILED ANALYSIS BEFORE THE RESULT:”
Method 2 Zero-Shot: “Provide only the final answer in the format below. Do not include any reasoning or explanation.” Chain-of-Thought: “Reason through this task step by step like a clinical expert, then return your final answer in the given format:”
Method 3 Zero-Shot: “Output your answer strictly using the following format. Explanations are not allowed.” Chain-of-Thought: “Let’s use step by step inductive reasoning, given the clinical nature of the task. Return your final answer in the following format:”
Method 4 Zero-Shot: “Respond using the exact format specified. Do not add any additional commentary.” Chain-of-Thought: “Determine the answer to the task given the clinical context in a step by step fashion. Give your final answer in the format below:”
Method 5 Zero-Shot: “Return the final answer exactly as outlined below without providing any justification or reasoning.” Chain-of-Thought: “Let’s break the task into multiple steps and reason step by step as a clinician would. Then, provide your final answer in the specified format:”
[…]
Model Prompt 1* Prompt 2 Prompt 3 Prompt 4 Prompt 5 Overall
Llama3.3-70B 3.10-3.10 (7.8%) 2.08-2.08 (5.28%) 2.75-2.75 (6.81%) 2.89-2.89 (7.25%) 2.24-2.24 (5.61%) 2.61-2.61 (6.55%)
Mistral-Large 3.40-3.40 (8.0%) 3.03-3.03 (7.14%) 3.35-3.35 (7.91%) 2.76-2.76 (6.64%) 2.99-2.99 (7.10%) 3.11-3.11 (7.36%)
Qwen2.5-72B 2.70-2.70 (6.5%) 2.24-2.24 (5.42%) 1.47-1.47 (3.54%) 3.27-3.27 (8.08%) 1.94-1.94 (4.73%) 2.32-2.32 (5.65%)
Athene-V2 2.40-2.40 (5.8%) 1.72-1.72 (4.16%) 1.19-1.19 (2.87%) 3.11-3.11 (7.63%) 1.70-1.70 (4.13%) 2.02-2.02 (4.92%)
* Prompt 1 is the original task instruction and also the one used in the main experiments.
A.5 LLM-as-a-Judge
To investigate LLM failure modes, we employed an LLM-as-a-judge framework implemented with OpenAI-o3 to annotate three binary metrics: hallucination, omission, and incompleteness. For each instance, the LLM judge provided a binary assessment indicating the presence of a failure mode, along with an explanatory statement citing specific examples or rationales. Validation against expert human annotations using the same criteria showed close alignment with the LLM-as-a-judge results.
Prompt of LLM-as-a-Judge for Error Analysis Your Task As an experienced medical professional with expertise in clinical documentation, diagnosis, and treatment planning, you will evaluate the accuracy and quality of multiple language model outputs on clinical text tasks. Assess responses for the specified error types and provide binary (0/1) judgments with a short, evidence-based justification for each model response. Input Components You will receive: Task Instructions: The clinical task the models were asked to perform, which describes the input information and the task requirements. Input: The text given to the models, which is sourced from patient records, clinical notes, or other relevant medical documentation. Answer: The correct answer for this instance, which serves as an important reference (i.e., what diagnosis or decision should be made based on the input). Model Responses: The responses generated by different language models, which should include the chain-of-thought (CoT) reasoning and a final answer, each with an anonymous identifier (e.g., Model_abc12345). Evaluation Criteria Evaluate each model response independently for: 1. Hallucination (Binary: 0 or 1) Definition: Any claim in the model response that is not supported by the provided Input or reflects incorrect/non-existent medical facts/knowledge. This includes: fabricated facts; invented numeric values (labs/vitals/doses); unsupported temporal/causal inferences; citing diagnoses/medications/history/guidelines absent from the Input; medically false statements (e.g., wrong contraindications, impossible dose ranges, pathophysiology that contradicts accepted evidence, misuse of units); or reliance on incorrect external knowledge. 1 = Hallucination: The response introduces any unsupported or medically false content—for example, a diagnosis not in the Input, a contraindication stated incorrectly, a guideline claim not evidenced by the Input, or incorrect medical knowledge. 0 = No Hallucination: All relevant claims are both (a) grounded in the Input (verbatim quote or faithful paraphrase) and (b) medically sound (consistent with established clinical knowledge). 2. Omission (Binary: 0 or 1) Definition: The model fails to include task-required, clinically relevant facts from the provided Input that are necessary to produce the Expected Output. This includes missing: key indicators for disease diagnosis, key qualifiers (onset, duration, timing, laterality), critical numeric details (dose/unit/frequency/range), severity/stage/grade, abnormal/critical flags, or fields explicitly mandated by the task/schema. 1 = Omission: One or more important and required information is ignored or incompletely reported so that the Expected Output is not fully met. This includes: missing the essential medical information, and over-generalization that drops required specificity (e.g., reporting ”antibiotic given” when Input specifies ”ceftriaxone 1 g q24h”). 0 = No Omission: All clinically significant and task-relevant information from the Input that is necessary to fulfill the Expected Output is present.
Clinical LLM Evaluation Prompt Continued 3. Incomplete (Binary: 0 or 1) Definition: The response is truncated or not fully executed, or its reasoning is insufficient to reach the conclusion (e.g., missing steps, unjustified leaps, internal inconsistencies). This includes: (i) output cut off or ending mid-sentence; (ii) required components/fields or mandated format/schema missing or only partially filled; (iii) reasoning gaps where key steps are skipped or the final answer conflicts with its own reasoning; (iv) use of unresolved placeholders (”TBD/…/—”) or vague generalities where specifics are required; (v) providing only a subset when the task requires exhaustive listing/coverage. 1 = Incomplete: The reasoning steps of model response are unfinished or incomplete. Any of the above conditions is present. 0 = Complete: The model’s reasoning steps is completed and response is finished successfully. Output Format Provide your evaluation as a JSON object with score and reason for each model. Specifically, the reason should explicitly show the evidence (e.g., raw text within Input) and point out the model errors. Use the model’s anonymous identifier as the key: { "Model_abc12345": { "hallucination": 0 or 1, "reason_of_hallucination": "several sentence explanation with examples or rationale.", "omission": 0 or 1, "reason_of_omission": "several sentence explanation with examples or rationale.", "incomplete": 0 or 1, "reason_of_incomplete": "several sentence explanation with examples or rationale." }, "Model_def67890": { "hallucination": 0 or 1, "reason_of_hallucination": "several sentence explanation with examples or rationale.", "omission": 0 or 1, "reason_of_omission": "several sentence explanation with examples or rationale.", "incomplete": 0 or 1, "reason_of_incomplete": "several sentence explanation with examples or rationale." }, ... } Evaluation Guidelines 1. Be Objective and Clinically Relevant: Your assessment should be grounded in medical knowledge and clinical practices and avoid subjective bias. 2. Concise Yet Justified Reason: Each reason should be clear, concise (several sentences), and medically sound, and explicitly show the evidence and score justification. 3. Ensure Precision: Ensure your evaluation is precise, consistent, and relevant to the specific clinical task requested. 4. Independent Evaluation: Evaluate each model independently, do not compare models to each other, only to the expected output and clinical standards. Now, please evaluate the provided model outputs using these criteria and return your evaluation in the specified JSON format with scores and reasons for models.